OpenStudy (anonymous):

Given the line -5x-2y=6 what is the slope-intercept form of the line, and the slope and y intercept?

OpenStudy (anonymous):

I know Y= 5x/2-3

OpenStudy (amistre64):

y = mx + b is the slope intercept form; where m=slope; and b = y int. it all amounts to solving for y thru algebra techniques

OpenStudy (amistre64):

-5x-2y=6 ; +5x -2y = 5x + 6 ; /-2 y = (-5/2)x -3

OpenStudy (amistre64):

slope = m = (-5/2) y int = b = -3

OpenStudy (anonymous):

I have no problem solving, I just can't seem to know what to do after that.

OpenStudy (anonymous):

thank you